# A systematic review of challenges and proposed solutions in modeling multimodal data

### v2

@@ 2.1.4 Imbalance in dimensionality
The term “imbalance in dimensionality” in the context of multimodal data fusion refers to situations in which the different modalities have significantly different numbers of dimensions or features. An imaging modality, for example, could include millions of pixels, but a genomic modality may have only a few hundred or thousand characteristics.
@@ 2.1.5 Optimal fusion technique
In addressing the challenge of finding the optimal fusion strategy, many studies in the literature explore different approaches. However, the choice can vary greatly depending on the specific problem. Ramachandram and Taylor 101 reviewed various options for learning an optimal architecture in the context of deep multimodal learning. They highlighted that determining the best fusion strategy depends on two architectural design choices. The first choice relates to the stage at which different modalities should be combined, which can be guided by considering the challenges and advantages of early, late, and intermediate fusion approaches. The second architectural design choice relates to which modalities to combine. They discussed the idea that including all available modalities may not always improve machine learning algorithm performance, suggesting the need for feature selection or optimization of fusion depth. They explored techniques for automatically learning the best ordering and depth of fusion during training to identify the most effective modalities to combine. These techniques include stochastic regularization, treating architecture optimization as a hyperparameter optimization problem using methods like Bayesian optimization, and employing incremental online reinforcement learning. A total of 13 studies in our review addressed this challenge. For example, Xu et al. 102 tackled this challenge by extending the neural architecture search 103, 104, which was originally designed for unimodal data, to incorporate the joint optimization of fusion strategies and modality-specific architectures for multimodal data. This extension allows for searching for unique architectures for distinct modalities and identifies the best strategies to fuse these architectures at the right representation level, and then they utilize state-of-the-art neural architecture search methods to explore and optimize fusion strategies for multimodal data. By jointly optimizing fusion strategies and modality-specific architectures, their approach aims to enhance prediction performance on EHR tasks while controlling for computation costs.
In the fusion of imaging and clinical data, Wang et al. 30 pointed out a common heuristic approach to data fusion in many studies, where data from multiple sources are simply concatenated without proper preprocessing or assessment of fusion quality. This failure to interpret the fused feature space can result in suboptimal model performance, particularly in medical imaging problems where the characteristics of different data sources can vary significantly. Therefore, they proposed modeling the fully connected layer of a deep neural network as a potential function following the classical Gibbs measure. By considering the features of the fully connected layer as random variables governed by state functions representing different data sources, the method quantifies source bias and aims to optimize data fusion quality. The use of a vector-growing encoding scheme, such as positional encoding, helps the conversion of low-dimensional clinical features into a rich feature space to complement high-dimensional imaging features, ultimately enhancing the fusion process in multi-source deep learning problems.
@@
2.2 Other challenges
Beyond the main challenges detailed in this manuscript, there is a range of unique additional issues that do not align with predefined categories yet are still critical for understanding and improving multimodal data integration. In LABEL:Tab:challenges, the specific challenges for “Other Challenges” are explicitly stated alongside their corresponding solutions, as they vary across different studies. One major challenge is extracting high-order interactions within and among data modalities, which is crucial for identifying complex, non-linear relationships that may exist between features. These interactions are particularly relevant in clinical contexts where modalities such as genomics and imaging provide complementary insights. Similarly, capturing the intricate relationships and temporal dynamics across modalities, while accounting for irregularities in patient visit timings, is a significant hurdle. Patient data often exhibits asynchronous patterns, with varying time gaps between visits, making it essential to develop models capable of handling such inconsistencies. High dimensionality adds another layer of complexity, as many modalities, such as imaging and genomics, have a vast number of features. This can lead to intensive computational requirements and the risk of overfitting, necessitating robust dimensionality reduction and feature selection techniques. Capturing intermodality interactions further complicates modeling, as it requires mechanisms that can integrate and prioritize information from heterogeneous sources without losing critical cross-modal correlations. Lastly, heterogeneity in data, whether in terms of formats, scales, or quality, presents a persistent challenge, particularly in ensuring that diverse modalities are integrated cohesively. These challenges highlight the need for innovative approaches, such as advanced attention mechanisms, hybrid fusion strategies, and efficient computational frameworks, to enable robust and interpretable multimodal data integration in clinical applications.
@@ 3 Discussion
In this systematic review of multimodal data fusion challenges and solutions in the clinical domains, we have identified and examined various obstacles faced when integrating data from diverse sources such as electronic health records, imaging, wearable devices, and more. Major obstacles were missing modality, limited datasets, interpretability concerns, imbalanced dimensionality, and selecting suitable fusion algorithms. The issue of small data continues to be a critical barrier for effective multimodal fusion, especially in medical settings where large datasets are often unavailable.
One notable aspect of this review is the variety of innovative approaches and methodologies proposed by researchers to address these challenges. For instance, methods such as multitasking learning, matrix completion, variational autoencoders, generative adversarial networks, hybrid graph convolutional networks, and knowledge distillation have been introduced to tackle missing of an entire modality and improve imputation accuracy. Additionally, techniques like data augmentation, transfer learning, and knowledge distillation have been effective in mitigating the impact of small datasets on model performance. However, their application in medical settings still remains constrained. Such methods often rely on access to high-quality pre-trained models or large, diverse datasets, which may not be readily available in the medical context. Furthermore, techniques like data augmentation and feature selection may struggle to fully capture the complex relationships between modalities when data is limited or heterogeneous. Interpretability, a critical aspect in medical applications, has been addressed through attention mechanisms, gradient-based attribution algorithms, and the incorporation of deep learning models with explanatory techniques like DeepSHAP. Moreover, strategies such as feature selection and dimensionality reduction have been proposed to combat the challenge of imbalance in dimensionality, ensuring fair representation across modalities. Determining the optimal fusion technique remains a key consideration, with early, intermediate, and late fusion strategies each offering advantages and drawbacks. Recent advancements in neural architecture search, weighted ensemble methods, and using attention mechanisms hold promise for optimizing fusion strategies and improving model performance.
This review was motivated by the potential of multimodal data modeling for chronic diseases such as COPD and asthma. However, we found a surprising lack of studies specifically addressing the challenges in integrating multimodal data for these conditions. Despite the increasing interest in leveraging diverse data sources, including omics, wearable devices, environmental data, and imaging, to improve the identification of endo- and phenotypes as well as disease progression classification, our systematic search yielded very few relevant studies. The few publications that did mention these diseases either did not focus on modeling challenges or met our exclusion criteria due to limited methodological insights. This gap highlights the need for future research to explore how multimodal integration can enhance disease management for COPD and asthma.
It is important to acknowledge certain limitations in this study. Firstly, the scope of the review may not encompass all existing methodologies and approaches in the field, and there may be additional strategies that were not included in our analysis. Moreover, the effectiveness of the solutions discussed herein may vary depending on the specific characteristics of the dataset and the complexity of the medical problem being addressed. Furthermore, while efforts have been made to select and present the most relevant research findings, there may be biases inherent in the selection process. Additionally, since the systematic search was conducted in October 2023, the rapidly evolving nature of the field means that new advancements may have emerged since the completion of this study.
@@
Future research should focus on specifically developing more tailored methods for handling small datasets in clinical contexts. This could involve incorporating domain-specific knowledge, improving data simulation techniques, or creating hybrid models that integrate both data-driven approaches and clinical insights. By overcoming the small data challenge, we can unlock the full potential of multimodal fusion models, enabling better decision-making and improving healthcare delivery across a variety of clinical applications.
4 Methods
4.1 Search strategy
To systematically identify relevant articles, a structured search strategy was employed (see Table 4). The search was conducted using two major databases, PubMed and manual search from Google Scholar. The time frame for the search was limited until October 2023 for PubMed (search date: 2023/10/09) and until end of 2023 for Google Scholar (search date: 2024/01/02). The strategy was designed to identify studies addressing challenges and methodologies in modeling multimodal data, with a focus on clinical applications. First, studies were required to explicitly mention “model fusion”, “data fusion,” or “multimodal” in their title, ensuring relevance to the topic of multimodal data modeling. To ensure comprehensive coverage, the Boolean operator OR was used within each keyword group to combine related terms. Next, the search included terms related to key research purposes, such as “disease prediction”, “diagnosis”, “clinical outcome prediction”, “modeling”, or “exacerbation”, to capture studies with clinical applications. Additionally, the search incorporated terms representing diverse data modalities, including “electronic health records”, “imaging data”, “wearable devices”, “longitudinal data”, “temporal data”, or “time-series data”, to ensure coverage of a wide range of multimodal datasets. To further refine the results, the term “challenges” was included to focus on studies discussing methodological or analytical obstacles related to multimodal data integration. Finally, the Boolean operator AND was used to combine these components. The final search query combined these elements to ensure a comprehensive yet focused selection of studies. In addition, a manual search was performed to identify any potentially relevant studies not captured in the database search.
4.2 Inclusion and exclusion criteria
4.3 Data extraction
For the data extraction process, we categorized the included publications according to the specific challenges they addressed. We identified five main technical challenges in modeling multimodal data: (i) missing entire modalities and incomplete datasets, (ii) small data (data sparsity), (iii) interpretability, (iv) imbalance in dimensionality, and (v) identifying the optimal fusion technique. We extracted information on the specific challenges each study tackled within the broader scope of the five identified challenges. We also captured the proposed solutions for overcoming these challenges. We further specifically collected information about the fusion strategies employed in each study. The choice of fusion technique directly relates to the challenge of finding the optimal fusion method (challenge v). Some strategies may also help mitigate other challenges, such as handling incomplete datasets (challenge i) or balancing dimensionality (challenge iv) by integrating information from multiple sources effectively. Finally, because different medical settings encounter distinct modeling challenges, we recorded the clinical domain of each study to link it with the challenges addressed. For example, missing modalities are particularly frequent in fields like neurology and oncology, reflecting the difficulty of integrating heterogeneous sources (such as imaging, genomics, and clinical notes) in these complex conditions Small datasets also occur more often in these fields because rare conditions or challenging data collection keep sample sizes low.
Five co-authors conducted the screening process. Selected full-text studies were distributed among MF, MW, MB, ChG, and JK for the eligibility check and extracting relevant data according to predefined criteria. After the independent evaluations, the results were cross-checked, and a final decision was made on which studies to include in the systematic review.
Data availability
@@ 4.3 Data extraction
All data supporting the findings of this study are available within the paper and its Supplementary Information files.
Acknowledgements
@@
Authors’ contributions
Competing interests
The authors declare that they have no competing interests.

### v3

@@ 2.1.4 Imbalance in dimensionality
The term “imbalance in dimensionality” in the context of multimodal data fusion refers to situations in which the different modalities have significantly different numbers of dimensions or features. An imaging modality, for example, could include millions of pixels, but a genomic modality may have only a few hundred or thousand features.
@@ 2.1.5 Optimal fusion technique
In addressing the challenge of finding the optimal fusion strategy, the problem in the context of deep multimodal learning can be reduced to two central architectural design choices. The first choice relates to the level at which different modalities should be combined, which can be guided by considering the challenges and advantages of early, late, and intermediate fusion approaches. The second architectural design choice concerns which modalities to combine, as including all available modalities does not always enhance the performance of a machine learning algorithm 101. A total of 13 studies in our review identified finding the optimal fusion strategy as a challenge. For example, Xu et al. 102 tackled this challenge by extending the neural architecture search 103, 104, which was originally designed for unimodal data, to incorporate the joint optimization of fusion strategies and modality-specific architectures for multimodal data. This extension allows for searching for unique architectures for distinct modalities and identifies the best strategies to fuse these architectures at the right representation level, and then they utilize state-of-the-art neural architecture search methods to explore and optimize fusion strategies for multimodal data. By jointly optimizing fusion strategies and modality-specific architectures, their approach aims to enhance prediction performance on EHR tasks while controlling for computation costs.
In the fusion of imaging and clinical data, Wang et al. 30 pointed out a common heuristic approach to data fusion in many studies, where data from multiple sources are simply concatenated without proper preprocessing or assessment of fusion quality. This failure to interpret the fused feature space can result in suboptimal model performance, particularly in medical imaging problems where the features of different data sources can vary significantly. Therefore, they proposed modeling the fully connected layer of a deep neural network as a potential function following the classical Gibbs measure. By considering the features of the fully connected layer as random variables governed by state functions representing different data sources, the method quantifies source bias and aims to optimize data fusion quality. The use of a vector-growing encoding scheme, such as positional encoding, helps the conversion of low-dimensional clinical features into a rich feature space to complement high-dimensional imaging features, ultimately enhancing the fusion process in multi-source deep learning problems.
@@
2.1.6 Other challenges
Beyond the main challenges detailed in this manuscript, there is a range of unique additional issues that do not align with predefined categories yet are still critical for understanding and improving multimodal data integration. In LABEL:Tab:challenges, the specific challenges for “Other Challenges” are explicitly stated alongside their corresponding solutions, as they vary across different studies. One major challenge is extracting high-order interactions within and among data modalities 55, 78, which is crucial for identifying complex, non-linear relationships that may exist between features. These interactions are particularly relevant in clinical contexts where modalities such as genomics and imaging provide complementary insights. Similarly, capturing the intricate relationships and temporal dynamics across modalities 53, 62, while accounting for irregularities in patient visit timings, is a significant hurdle. Patient data often exhibits temporal patterns, with varying time gaps between visits, making it essential to develop models capable of handling such inconsistencies. High dimensionality adds another layer of complexity 54, 65 as many modalities, such as imaging and genomics, have a vast number of features. This can lead to intensive computational requirements and the risk of overfitting, necessitating robust dimensionality reduction and feature selection techniques. Capturing intermodality interactions 19, 21, 37, 53, 58, 70, 73 further complicates modeling, as it requires mechanisms that can integrate and prioritize information from heterogeneous sources without losing critical cross-modal correlations. Lastly, heterogeneity in data 42, 43, 52, 55, 63, 76, 80, whether in terms of formats, scales, or quality, presents a persistent challenge, particularly in ensuring that diverse modalities are integrated cohesively. These challenges highlight the need for innovative approaches, such as advanced attention mechanisms, hybrid fusion strategies, and efficient computational frameworks, to enable robust and interpretable multimodal data integration in clinical applications.
@@ 3 Discussion
In this systematic review, we identified several key technical challenges that hinder the modeling of multimodal medical data and assessed how recent studies have attempted to overcome them. The most commonly addressed obstacles were missing modalities and small sample sizes, which were tackled in 15 (22%) and 17 (25%) of the 69 studies, respectively.
A notable aspect of this review is the variety of innovative solutions proposed to address the key challenges in multimodal modeling. For missing data modalities, researchers have introduced methods such as multitask learning 43, matrix completion 79, 87, variational autoencoders 23, 90, generative adversarial networks 51, 71, hybrid graph convolutional networks 20, and knowledge distillation wang_multi-modal_2023. These approaches aim to impute or compensate for absent modalities and have shown improvements in handling incomplete datasets. To overcome limited sample sizes (“small data”), strategies such as data augmentation 25, 15, transfer learning from large pretrained models 29, 22, 15, 41, 59, 50, and knowledge distillation 56 have been employed to maximize the value of scarce data. However, the effectiveness of these techniques in clinical contexts is constrained: many rely on external pretrained models or large datasets that are not always available, and they may fail to capture complex cross-modal patterns when data are sparse or heterogeneous. For improving interpretability, studies have applied attention mechanisms, gradient-based attribution tools 21, 22 , and explanation techniques such as DeepSHAP 42, 44, xu_explainable_2021, 46, which help clarify model decision-making. To address imbalances in the dimensionality of different modalities, where one modality may dominate due to feature richness, methods like feature selection, dimensionality reduction 80, and weighted loss functions 39, 25, 58, 22 have been used to achieve more balanced integration. Finally, the selection of an optimal fusion strategy remains an open challenge. Early, intermediate, late, and hybrid fusion each offer distinct strengths and weaknesses, and recent advances such as neural architecture search and attention-based fusion 19, 26, 45 show promise in tailoring integration methods to specific tasks. Taken together, these findings highlight that there is no one-size-fits-all solution: each challenge requires targeted techniques, and many multimodal pipelines combine multiple methods to achieve robust performance.
While summarizing these approaches, it is important to clarify that this review does not introduce new methods or propose original solutions; rather, it systematically synthesizes techniques already published in the literature. Earlier reviews have examined multimodal machine learning in narrower contexts, for example, focusing on fusing imaging modalities, integrating omics data, or combining clinical data like EHRs with images, and have typically been organized by data type or algorithmic technique. By contrast, we systematically catalogued methods according to the specific modeling challenge they intend to solve. This approach allows us to directly map each problem (missing data, small data, etc.) to the repertoire of solutions that have been proposed, thus highlighting which areas have many established methods and which remain relatively underexplored. To our knowledge, this is the first review to systematically catalogue current methods by the specific modeling challenge they address across diverse multimodal data types in medical applications.
By organizing the literature according to challenges, we make it easier for researchers and practitioners to pinpoint relevant methods when faced with a particular problem. For example, if a new project is struggling with a missing data modality or an imbalance between imaging and tabular data, our compiled findings serve as a reference to quickly identify candidate solutions (such as data imputation techniques or feature-balancing approaches) that have been tried previously. Moreover, our review highlights which challenges have a rich range of methods and which do not. We observed, for instance, that numerous techniques exist to address missing modalities (ranging from traditional imputation to sophisticated cross-modal generative models), whereas approaches to ensure interpretability of multimodal models are relatively scarce in the literature. Given the critical importance of model transparency in healthcare settings, this imbalance suggests that future research needs to more directly tackle interpretability, for example, by developing multimodal-specific explainability tools or evaluation frameworks for model explanations. Similarly, we noted that the small data challenge, although frequently mentioned, remains difficult to overcome; even the most common approaches (augmentation and transfer learning) have inherent limitations when data are extremely limited or not representative. By illuminating these gaps, our work contributes to the community’s understanding of where additional methodological innovation or investment of effort is most needed. Ultimately, the insights from our review help map the landscape of multimodal data fusion in medicine: they show where the field has made significant progress (e.g. handling missing information) and where key opportunities for improvement lie (such as improving model interpretability and developing techniques that do not depend on large datasets).
It is important to acknowledge certain limitations in this study. First, our review may not encompass all existing methodologies in this rapidly evolving field; some relevant strategies might not have been captured by our search or inclusion criteria. Second, the effectiveness of the solutions we discuss can be context-dependent; what works for one clinical dataset or problem might not generalize to others, so these solutions should be applied with consideration of specific data characteristics. Third, although we aimed for comprehensive coverage, it’s possible that certain studies or viewpoints were missed. Fourth, because our systematic search was last conducted in October 2023, the field has likely advanced since then and new techniques or findings may have emerged post-search that are not included in our review. Fifth, it is worth noting that although we initially highlighted chronic diseases such as COPD and asthma as motivating example due to their multimodal nature involving imaging, wearable, environmental, and clinical data, our systematic search yielded very few studies explicitly focused on these conditions. This gap may reflect a true underrepresentation of such diseases in current multimodal modeling research, rather than a failure of the search strategy itself, as our queries were designed to capture a broad range of clinical domains. We acknowledge this limitation and highlight it as an important area for future research. Finally, we would like to point out that although our search was not limited to deep learning-based approaches, the majority of included studies use deep learning techniques. The review is therefore focused on deep learning methods by result, rather than by design.
Beyond methodological issues, broader concerns of fairness, ethics, and deployment are increasingly recognized as critical in multimodal AI for medicine. Recent reviews emphasize that fairness research is still scarce in many clinical domains, with sensitive attributes such as race, sex, and socio-economic status often underrepresented in available datasets, thereby limiting equity in model performance liu2023translational, norori2023fairness. In medical imaging, for example, studies have shown that algorithms can exploit demographic shortcuts that yield good results on benchmark datasets but fail to generalize to new populations yang2024limits. The multimodal setting introduces additional complexity, as different modalities may carry distinct sources of bias, which can interact or even reinforce each other hickmon2024multimodal. Beyond fairness, deployment in clinical practice raises ethical and practical concerns including data privacy, informed consent for the integration of heterogeneous data sources, interpretability of multimodal decisions for clinicians, and alignment with regulatory frameworks. While addressing these aspects was beyond the scope of our systematic review, they represent important directions for future work to ensure that technical advances in multimodal modeling translate into safe, equitable, and clinically useful applications.
@@
Future research should focus on developing more tailored methods for handling small datasets in clinical contexts. This could involve incorporating domain-specific knowledge, improving data simulation techniques, or creating hybrid models that integrate both data-driven approaches and clinical insights. By overcoming the small data challenge, we can unlock the full potential of multimodal fusion models, enabling better decision-making and improving healthcare delivery across a variety of clinical applications.
4 Conclusion
In this systematic review, we synthesized evidence from 69 studies and found that multimodal data fusion holds significant promise in clinical domains but that several challenges persist, particularly regarding missing modalities, dimensionality imbalance, interpretability, and optimal fusion techniques. Among these, the issue of small data continues to be a critical barrier for effective multimodal fusion, especially in medical settings where large datasets are often unavailable. Although various strategies (e.g., multimodal pre-training, transfer learning, data augmentation) have been proposed to address the small-data issue, they often depend on resources like large pre-trained models or extensive datasets that are not readily available in medical contexts, limiting their practical impact so far. Furthermore, techniques like data augmentation and feature selection may struggle to fully capture the complex relationships between modalities when data is limited or heterogeneous. Our review emphasizes that while these approaches show promise, their real-world applicability in medical settings is limited, requiring further refinement to handle the unique challenges of clinical data. Addressing the small data issue is crucial for maximizing the effectiveness of multimodal fusion models and making them more robust and generalizable across various healthcare scenarios. Future research should focus on developing tailored methods to effectively handle the challenge of small datasets in clinical contexts. This could include approaches like integrating rich domain knowledge into modeling, improving data simulation/augmentation techniques, or designing hybrid models that combine data-driven algorithms with clinical insight. By overcoming the small-data challenge, the field can unlock the full potential of multimodal fusion models, enabling more informed decision-making and improved healthcare delivery across a variety of clinical applications.
Data Availability Statement
@@ 4 Conclusion
All data supporting the findings of this study are available within the paper and its Supplementary Material.
Author Contributions
Funding
@@
Acknowledgments
We thank Daniela Karakuleva for her assistance with the extraction of data modalities and performance metrics.
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Generative AI statement
The authors declare that no Gen AI was used in the creation of this manuscript.
Supplemental Material
The Supplementary Material for this article includes: Supplementary Table S1. Extracted data from the 69 included studies, provided as a separate Excel file. The table contains information on clinical domains, fusion strategies, data modalities, challenges and corresponding solutions, task types, and performance metrics. The Supplementary Material can be found online at: